OpenStudy (anonymous):

I would absolutely love and adore some help integrating (6e^6t)/(e^(12t) +12 e^(6t) +27) dx

OpenStudy (anonymous):

\[ I=\int\frac{6e^{6t}}{e^{12t}+12e^{6t}+27}dt \] start with the substitution \[e^{6t}=u\]

OpenStudy (anonymous):

so I should get to: \[I=\int\limits_\ \frac{u}{u^{2}+12u+27} du\] correct?

OpenStudy (anonymous):

which doesn't look scary at all

OpenStudy (anonymous):

yes. factorize the denominator

OpenStudy (anonymous):

Here's where I'm at so far:\[=6 \int\limits_{} \frac{u}{(u+9)(u+3)} du\] \[\therefore\frac{u}{(u+9)(u+3)}=\frac{A}{u+9}+\frac{B}{u+3}\] \[\therefore u=A(u+3)+B(u+9)\]

OpenStudy (anonymous):

set u=-9 \(-9=A(-9+3)\implies A={3\over2}\) set a=-3 \(-3=C(-3+9)\implies B=-{1\over2}\)